Find the requested function and state the domain
find g 0 f, for f(x)=6x-5 and g(x)=6x+6

Answer by stanbon(75887)

Find the requested function and state the domain
find g 0 f, for f(x)=6x-5 and g(x)=6x+6
----------
gof = g[f(x]] = g[(6x-5)] = 6(6x-5)+6 = 36x-30+6 = 36x-24
============================================================
Cheers,
Stan H.